ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 234, 2019

Akagi, G. Fractional flows driven by subdifferentials in Hilbert spaces  809
Akofor, E. On Lipschitz retraction of finite subsets of normed spaces  777

Andrews, U., G. Igusa, J. S. Miller and M. I. Soskova Characterizing the continuous degrees

743

Araújo, J., P. J. Cameron, C. Casolo and F. Matucci Integrals of groups

149

Baltazar, H., R. Diógenes and E. Ribeiro Jr. Isoperimetric inequality and Weitzenböck type formula for critical metrics of the volume

309

Cameron, P. J. See J. Araújo Casolo, C. See J. Araújo Cholak, P. A., G. Igusa, L. Patey, M. I. Soskova and

D. Turetsky The Rado path decomposition theorem

179

Cohen, S. and S. Shelah Generalizing random real forcing for inaccessible cardinals

547

Cohen, A. M. and D. B. Wales Corrigendum to “Brauer algebras of simply laced type”

499

Cui, Y. See S. Shang Diógenes, R. See H. Baltazar van Gool, S. J. and B. Steinberg Pro-aperiodic monoids via

saturated models

451

Harpaz, Y., J. Nuiten and M. Prasma Tangent categories of algebras over operads

691

Hooda, N. and M. Thakur Rational subgroups and invariants of F_{4}  581
Huang, H. John’s position is not good for approximation  37

Igusa, G. See P. A. Cholak Igusa, G. See U. Andrews Isaacs, I. M., G. Navarro and B. Sambale Irreducible extensions

of characters

629

Jespers, E., D. Riley and M. Shahada On polynomials that are not quite an identity on an associative algebra

371

Kamber, A. L^{p}-Expander Graphs  863

Kopeliovich, Y. and S. Zemel On spaces associated with invariant divisors on Galois covers of Riemann surfaces and their applications 393 Kuhlmann, F.-V. Elimination of ramification II: Henselian

rationality

927

Leiderman, A. G., S. A. Morris and M. G. Tkachenko The separable quotient problem for topological groups

331

Martinsson, A., F. Mousset, A. Noever and M. Trujić The edge-statistics conjecture for ℓ ≪ k^{6/5}

677

Matucci, F. See J. Araújo Melnikov, A. G. and K. M. Ng The back-and-forth method and

computability without delay

959

Meunier, F. and S. Zerbib Envy-free cake division without assuming the players prefer nonempty pieces

907

Miller, J. S. See U. Andrews Morris, S. A. See A. G. Leiderman Mousset, F. See A. Martinsson Navarro, G. See I. M. Isaacs Newman, I. I. and Y. Rabinovich On connectivity of the facet

graphs of simplicial complexes

521

Ng, K. M. See A. G. Melnikov Noever, A. See A. Martinsson Nuiten, J. See Y. Harpaz Patey, L. See P. A. Cholak

Patrão, M. The topological entropy of endomorphisms of Lie groups  55

Pollack, A., C. Wan and M. Zydor A G_{2}-period of a Fourier coefficient of an Eisenstein series on E_{6}

229

Prasma, M. See Y. Harpaz

Preusser, R. The V -monoid of a weighted Leavitt path algebra  125

Rabinovich, Y. See I. I. Newman Ramras, D. A. The homotopy groups of a homotopy group

completion

81

Ribeiro, E., Jr. See H. Baltazar Riley, D. See E. Jespers Sambale, B. See I. M. Isaacs

Schneider, F. M. Equivariant dissipation in non-archimedean groups  281

Shahada, M. See E. Jespers Shang, S. and Y. Cui Continuity points and continuous selections of

the set-valued metric generalized inverse in Banach spaces

209

Shelah, S. See S. Cohen

Solan, O. N. Stable and well-rounded lattices in diagonal orbits  501

Soskova, M. I. See U. Andrews Soskova, M. I. See P. A. Cholak

Srinivasan, P. A virtually ample field that is not ample  769

Steinberg, B. See S. J. van Gool Thakur, M. See N. Hooda Tkachenko, M. G. See A. G. Leiderman Tringali, S. Structural properties of subadditive families with

applications to factorization theory

1

Trujić, M. See A. Martinsson Turetsky, D. See P. A. Cholak Wales, D. B. See A. M. Cohen Wan, C. See A. Pollack Zemel, S. See Y. Kopeliovich Zerbib, S. See F. Meunier Zühlke, P. Homotopical and topological rigidity of hypersurfaces of

spherical space forms

645

Zydor, M. See A. Pollack